OpenStudy (314dokg):

how do you add 3 of the same trgio functions? example: sin 0 + sin 90 + sin 180 i did the formula sin(a) + sin(b) = (2)(sin(a+b/2)(cos(a-b/2) but its gud for 2 only... help me for 3

OpenStudy (314dokg):

to make it easier to understand \[\sin (a) + \sin(b) = 2 \times \sin(\frac{ a+b }{ 2 }) \times \cos(\frac{ a-b }{ 2 })\]

ganeshie8 (ganeshie8):

why to do all that circus ? dont we knw that sin 0 = 0, sin 90 = 1, sin 180 = 0 ?

ganeshie8 (ganeshie8):

subbing them and adding gives, 0 + 1 + 0 = 1

ganeshie8 (ganeshie8):

??

OpenStudy (314dokg):

sorry. we just started trigo. its our assignment and i hav no clue.

ganeshie8 (ganeshie8):

ohk,

ganeshie8 (ganeshie8):

did they cover unit circle yet ?

ganeshie8 (ganeshie8):

did they teach u about unti circle ?

OpenStudy (314dokg):

yep i hav the table... but we hav to evaluate it. and show the proccess

OpenStudy (314dokg):

i knw their value in the quadrantal angles and stuff. all i need is the process

ganeshie8 (ganeshie8):

ok i wont be useful i guess, im not getting the purpose of q oly :|

OpenStudy (314dokg):

q?

ganeshie8 (ganeshie8):

question

OpenStudy (314dokg):

oh ok then... but ur first reply... all you hav to do is to replace the value of sine. like sin 0 = 0 and then just add 3 of then so we'll get 1? is that the easy way?

ganeshie8 (ganeshie8):

i wud do that lol, the other way u trying makes no sense to meh

OpenStudy (314dokg):

oh i see. nyways thnx alot :)

ganeshie8 (ganeshie8):

np :)